a rectangle has a perimeter of 96 cm .its length is twice its breadth. find its dimensions.
Answer by oberobic(2304)

p = 2(l+w), where p=perimeter, l=length, and w=width.
p = 96 cm, given
l = 2*w, given
96 = 2(2*w + w) = 4w + 2w = 6w
6w = 96
w = 16
l = 2*w = 32
.
check the perimeter
2*(16+32) = 2*48 = 96
correct.
.
Answer: length = 32 cm and width = 16 cm
.
Done.